A 7.0 Ohms resistor is connected in the series with another resistor and a 4.5 V battery. The current is 0.60 A. calculate the value of the unknown resistance.




1
Expert's answer
2021-06-03T18:26:32-0400

"V=I(r+r_x)\to4.5=0.6\cdot7+0.6\cdot r_x\to r_x=0.5\ (\Omega)" . Answer